# Competing color superconductivity and color Kondo effect in quark matter

### v1

@@ Appendix C Bethe Ansatz equations
To diagonalize the transfer matrix, we use the nested $T-Q$ relations and functional Bethe Ansatz method. Note that since the rank of $\mathfrak{su}(3)$ is 2, there are two ranks of Bethe Ansatz equations for the color space and hence there are two sets of Bethe roots and two Baxter $Q$-functions defined by
@@
To construct the transfer matrix we define the forward and reflected monodromy matrices acting on the auxiliary space $A\cong\mathbb{C}^{3}$ as
@@
$[\tau(u),\tau(v)]=0\quad\forall\,u,v\in\mathbb{C}.$ (86)
@@ D.1 Kondo phase
[…]
@@ D.1.1 Fundamental excitations
[…]
@@
The highest-weight state and its nine descendants are plotted in the $Y$$I_{3}$ plane in Fig. 6.
@@
These excitations carry Dynkin labels $(p,q)=(0,3)$, which is 10-dimensional, and the highest weight state carries $Y_{\rm max}=2$ and $I_{3}=0$. The highest weight state and the nine descendant states are plotted in the $Y$$I_{3}$ plane in Fig. 7.
@@ D.2 The YSR phase
The construction of this state is different when $\frac{1}{2}<\delta<1$ and when $21<\delta<$ 2. In the former case, there exists a unique solution of Bethe Ansatz equations of rank-1 of the form
@@
Adding this solution, we obtain (for parameter range $11/2<\delta<$ 1), both rank-1 and rank-2 equation changes as follows:
@@
Before we compute the energy of the bound mode, we now turn to the construction of this state for the parametric range $21<\delta<$ 2. In this regime, the rank-2 equation given by Eq. (166) becomes
@@
All the excitations constructed in Sec.D.1.1 are valid excitations in the bulk in this regime also. And the above-constructed boundary excitations are not all possible boundary expectations. For example, one unique triality 2 boundary excitation is possible in the parametric regime $21<\delta<3/$ 2 by adding only the boundary string in the rank-1 Bethe equations such that the Bethe Ansatz equations in Fourier space become
@@ D.3 The Unscreened regime
[…]
@@
The total number of rank-1 roots including the one purely imaginary boundary string solution is $M_{1}=1+\int\rho_{1}(\lambda)\mathrm{d}\lambda=\frac{1}{3}\left(2N_{q}+3\right)$ and the total number of rank-2 roots including the one purely imaginary boundary string solution is $M_{2}=1+\int\rho_{2}(\mu)\mathrm{d}\mu=\frac{1}{3}\left(N_{q}+3\right)$. Since the Dynkin labels $(p,q)=(0,-1)$, this is not a valid state as one of the Dynkin labels is negative. Thus, to make it a valid state, we need to add a hole in the rank 2 equations at position $\vartheta$, such that the new Bethe Ansatz equations become

### v2

@@ Appendix C Bethe Ansatz equations
To diagonalize the transfer matrix, we use the nested TQT-Q relations and functional Bethe Ansatz method. Note that since the rank of $\mathfrak{su}(3)$ is 2, there are two ranks of Bethe Ansatz equations for the color space and hence there are two sets of Bethe roots and two Baxter $Q$-functions defined by
@@
To construct the transfer matrix, we define the forward and reflected monodromy matrices acting on the auxiliary space $A\cong\mathbb{C}^{3}$ as
@@
$\tau(u),\tau(v)]=0\quad\forall\,u,v\in\mathbb{C}.$ (86)
@@ D.1 Kondo phase
$\sum_{\beta=1}^{M_{1}}F(\lambda_{\beta})\to\int_{-\infty}^{\infty}F(\lambda^{\prime})\rho_{1}(\lambda^{\prime})d\lambda^{\prime},\quad\sum_{k=1}^{M_{2}}G(\mu_{k})\to\int_{-\infty}^{\infty}G(\mu^{\prime})\rho_{2}(\mu^{\prime})d\mu^{\prime}.$
@@ D.1.1 Fundamental excitations
$E_{\overline{\lambda}}=D\tan^{-1}\left(\frac{2b}{\overline{\lambda}^{2}-b^{2}+1}\right)-iD\int\frac{e^{-|\omega|-ib\omega}\cos\left(\omega\overline{\lambda}\right)}{\omega}d\omega=0$ (129)
@@
$I_{3}$YY(32,1)($\frac${3}{2},1)(12,1)($\frac${1}{2},1)(12,1)(-$\frac${1}{2},1)(32,1)(-$\frac${3}{2},1)(1,0)(1,0)(0,0)(0,0)(1,0)(-1,0)(12,1)($\frac${1}{2},-1)(12,1)(-$\frac${1}{2},-1)(0,2)(0,-2)
@@
$I_{3}$YY(0,2)(0,2)(12,1)($\frac${1}{2},1)(12,1)(-$\frac${1}{2},1)(1,0)(1,0)(0,0)(0,0)(1,0)(-1,0)(32,1)($\frac${3}{2},-$1)(\frac{1}{2},$-$1)(-\frac{1}{2},$-$1)(-\frac{3}{2},$-1)
@@ D.2 The YSR phase
The construction of this state is different when $\frac{1}{2}<\delta<1$ and when $1<\delta<2$. In the former case, there exists a unique solution of Bethe Ansatz equations of rank-1 of the form
@@
Adding this solution, we obtain (for parameter range $1/2<\delta<1$), both rank-1 and rank-2 equation changes as follows:
@@
Before we compute the energy of the bound mode, we now turn to the construction of this state for the parametric range $1<\delta<2$. In this regime, the rank-2 equation given by Eq. (166) becomes
@@
All the excitations constructed in Sec.D.1.1 are valid excitations in the bulk in this regime also. And the above-constructed boundary excitations are not all possible boundary expectations. For example, one unique triality 2 boundary excitation is possible in the parametric regime $1<\delta<3/2$ by adding only the boundary string in the rank-1 Bethe equations such that the Bethe Ansatz equations in Fourier space become
@@ D.3 The Unscreened regime
$M_{1}=\int d\lambda\tilde{\rho}_{1}(\lambda)=\frac{2N_{q}}{3},\qquad M_{2}=\int d\mu\tilde{\rho}_{2}(\mu)=\frac{N_{q}}{3}.$
@@
The total number of rank-1 roots including the one purely imaginary boundary string solution is $M_{1}=1+\int\rho_{1}(\lambda)\mathrm{d}\lambda=\frac{1}{3}\left(2N_{q}+3\right)$ and the total number of rank-2 roots including the one purely imaginary boundary string solution is $M_{2}=1+\int\rho_{2}(\mu)\mathrm{d}\mu=\frac{1}{3}\left(N_{q}+3\right)$. Since the Dynkin labels (p,q)=(0,1)(p,q)=(0,-1), this is not a valid state as one of the Dynkin labels is negative. Thus, to make it a valid state, we need to add a hole in the rank 2 equations at position $\vartheta$, such that the new Bethe Ansatz equations become